Find the value(s) of x such that the angle between the vectors (0, 1, −1) and (−1, x, 0) is 2π/3 . Show all your calculations


1
Expert's answer
2021-07-16T01:59:46-0400

Solution

We know ,if x and y are two vectors ,the angle "\theta" between them is given by;

Cos "\theta" ="\frac{(\vec{x}.\vec{y})}{\Vert x\Vert\vert y\Vert}"

Cos("\frac{2π}{3}")=-"\frac12"

"\vec{x}.\vec{y}" =0(-1)+x (1)+0(-1)=x

"\Vert \vec{x}\Vert" ="\sqrt{0^2+1^2+(-1)^2}" ="\sqrt{2}"

"\Vert\vec{y}\Vert" ="\sqrt{(-1)^2+(x)^2+0^2}" ="\sqrt{1+x^2}"

We now equate;

"-\frac 12" ="\frac{x}{\sqrt{2(1+x^2)}}"

2(1+x2)=4x2

x2=1

x="\sqrt{1}" =1 or -1

If x =1

Cos "\theta" ="\frac 12" "\neq" Cos"(\frac{2π}{3})"

Chose X=-1 that gives the cos"\theta" as "-\frac12"

Answer

x=-1